# skipabove=0.5skipbelow=0.5align=center A Research Agenda to Democratize Unit Proof Memory Safety Verification A Research Agenda for Unit Proof Verification

### v1

@@ Abstract
Software underlies digital and cyber-physical systems, including industrial, medical, financial, and aviation infrastructure [1]. Code-level defects, such as arithmetic errors and memory corruption vulnerabilities, compromise system safety and security and cause significant losses [2]. For example, a null-pointer bug in CrowdStrike Falcon in 2024 impacted global transportation systems and caused over $5 billion in damages [3]. Traditional engineering practices, such as systematic or randomized testing [4, 5], program analysis [6, 7], and runtime sandboxing[8], are commonly used to detect or mitigate such defects. However, these methods validate software only approximately and cannot guarantee correct behavior on all inputs. In contrast, formal verification [9] proves software correctness, guaranteeing no code-level defects. This raises an important (and perennial) question [10]: how to make formal verification more cost-effective for software engineers?
Engineering tools and processes help reduce the cost of applying a technology. For formal methods, there are code-level verification tools, e.g., the C and Rust bounded model checkers (CBMC[11], Kani [12]). To improve both engineering and tool scalability, engineers decompose the software and verify individual functions within it via unit proofs [13]. These unit proofs model the function’s interaction with other software components and verify the function using this model. We refer to the process of developing, using, and maintaining unit proofs as unit proofing. Organizations such as AWS [14], ARM [15], and the Linux Foundation [16] have adopted unit proofing to ensure software correctness and memory safety.
@@ II-A Formally Verifying Software Implementations
Many tools can verify low-level software implementations using a variety of formal methods [18, 19, 11]. This paper focuses on the subset of verification tools that use Bounded Model Checking (BMC) [20] as they require the least human effort [21], can automatically verify properties like memory safety, and are being adopted by major software organizations [14, 15]. We refer the reader to other works for comparisons of such tools [22, 23, 21].
@@
BMC has been applied in both industry (e.g., AWS boot firmware [25], ARM confidential computing firmware [15]) and open-source projects (e.g., Firecracker, Hifitime, S2N-Quic [26]). Beyer et al. [27] report that BMC tools find significantly more bugs in less time compared to testing and fuzzing tools. However, BMC struggles with state space explosions when applied to large software systems.
@@ II-B Unit Proofing for Enabling Code-level Verification
To apply BMC tools in practice, several papers have described similar engineering processes [14, 15, 16]. One example is AWS’s verification of embedded software and data structure libraries [14]. In their process, engineers develop unit proofs and use them to verify individual functions in the software. We refer to this process as “Unit Proofing”.111Chong et al. [14] did not name the process they described. We selected this name due to similarities between the described process and unit testing. As Figure 1 shows, these unit proofs model the function’s environment, including input, resources, and possibly other functions. However, the processes described in these papers require engineers to manually develop the unit proof models or provide their specifications, which can be costly and error prone.
Unit proofing shares similarities with unit testing and compositional verification, but is substantially distinct. Like unit tests, AWS’s unit proofs are developed together with the software, stored in the same repository, and executed in continuous integration workflows [14]. However, unit tests check correctness on specific inputs, while unit proofs verify correctness across all possible inputs within the model [28]. Compared with compositional verification, unit proofing is an engineering process that enables compositional verification at the code level. Traditional CV techniques [29, 30] only operate on abstract software models.
@@ III Exploratory Study of Unit Proofing
We evaluated this question using FreeRTOS unit proofs and reported vulnerabilities (CVEs). FreeRTOS was chosen due to its widespread use and available unit proofs [32, 33]. CVEs were used as they typically reflect critical memory safety vulnerabilities, which the unit proofs intend to address [33].
@@ III-A Methodology
Vulnerability Recreation: Following the vulnerability description, we identified and removed any input validations in the function that fixed the vulnerability.
@@ III-B Results
Only 5/11 recreated vulnerabilities were exposed by FreeRTOS unit proofs (Table I). With changes to the unit proofs, all vulnerabilities were exposed and one new high-severity vulnerability found (CVE-2024-BLINDED). Unit proofs are effective, but unit proof engineers can err.
@@
| CVE-2018-16527 | No | Insuff. Input Modeling | Development |
| CVE-2018-16600 | No | Insuff. Input Modeling | Development |
| CVE-2018-16603 | No | Insuff. Input Modeling | Development |
@@ III-C1 Insufficient Verification Properties (CVE-2018-16523)
BMC tools can verify program-agnostic properties, but require the properties to be enabled. The unit proof for CVE-2018-16523, a division-by-zero defect, missed the issue because the division-by-zero property was disabled 222From CBMC v6, this property is enabled by default.. The rationale behind disabling specific correctness checks remains unclear.
@@
Unit proofs must model inputs, broadly enough to cover all call sites yet precisely enough to avoid false positives. The CVE-2018-16527 vulnerability only occurs when the header of the input packet is truncated. The unit proof for the affected function incorrectly assumed the input packet was always large enough to contain the header.
@@ III-C4 Out-of-date Unit Proofs (CVE-2018-16525)
The vulnerable function’s signature was changed but the signature of the function called in the unit proof was not updated. As CBMC supported making calls to undefined functions and only reported coverage on the reachable lines of code, the updated function was no longer verified. When we fixed this unit proof, we exposed and reported a new vulnerability 333In response to our disclosure, AWS’s engineers changed CBMC’s default behavior for undefined functions, reporting errors instead of warnings. This led to many changes across their CBMC-using projects..
@@ IV Research Agenda: A Unit Proofing Framework
IV-A Unit Proofing Challenges and Open Questions
Here, we discuss the goals of each unit proofing stage and identify Open Questions (OQs) necessary to solving them.
@@ IV-A Unit Proofing Challenges and Open Questions
Unit proofing involves decomposing software into function units and verifying each one. The decomposition strategy impacts the number of unit proofs, the complexity of their models, and verification tractability. Smaller units are solvable but require increased unit proofing effort as more unit proofs with more complex models need to be created. Larger units have reduced unit proofing effort but may result in units that are too large for existing constraint solvers. Hence, an optimal decomposition should minimize unit proofing efforts but ensure all units are solvable.
@@
Software evolves, and section III shows that failing to update unit proofs leads to incorrect verification. Since unit proofs verify interacting functions, changes in one proof can invalidate others. Frequent software changes can make it impractical to regenerate or validate unit proofs after every change. As Chong et al. [14] note, reverifying software after each change is also costly. Understanding how software changes will affect unit proofs, and repairing misaligned unit proofs, will ensure unit proofs remain consistent with the software.
@@ IV-B1 Software Decomposer
The software decomposer will use novel algorithms to optimize decomposition by minimizing the number of units while ensuring each unit’s verification completes within a set time. This requires new methods to estimate model complexity and verification time. Building on prior research linking program features to specification size and effort [36, 37, 38, 34], we will study unit proofs to understand how program features impact model complexity and verification duration.
Insights from the study will inform a cost estimator that analyzes function features to estimate complexity and verification time. The decomposer algorithm will start from entry points and decompose functions if their estimated cost exceeds a threshold, iteratively replacing the most expensive function with its model. We will also explore a bottom-up design using graph decomposition algorithms [39, 40, 41, 42], partitioning the call graph by weighing nodes and edges based on verification time and modeling complexity to minimize total costs.
@@ IV-B2 Proof Builder
A unit proof defines models for inputs, shared resources, and called functions, requiring sound and complete models for accurate verification. In stateful systems or those using shared resources, the resource model depends on prior execution sequences. Traditional learning-based assume-guarantee reasoning [43, 44] refines assumptions iteratively using counterexamples, but applying this to functions with multiple interactions is costly due to the need for repeated model checking [45].
To address this, we propose to use a property-violation guided learning approach, starting with generic proof models and refining the models only until all reachable lines of code are covered and violations reported by the verification tool are resolved. Refinements will be informed by interventions from the proof repairer (section IV-B3). The learned models will then be propagated as assertions to the respective components, with violations indicating defects. This approach will also help indicate specific execution paths that trigger a defect.
@@ IV-B3 Proof Repairer
The proposed proof repairer will detect faults in unit proofs and propose intelligent fixes. To design the repairer, we will first conduct an empirical study on existing and new unit proofs, assess their coverage, violation reports, and ability to expose artificial bugs. We will iteratively resolve identified faults and use records of the necessary fixes to build a taxonomy of faults and interventions.
Using insights from this taxonomy, the proof repairer will identify and resolve faults in unit proofs. Since each program is unique, fixing faults requires a deep understanding of program semantics. Traditional program analysis techniques are fast but imprecise, while large language models (LLMs) can reason about code [46, 47] but are slower and costly. We will explore two hybrid approaches: one using search-based repairs with LLMs as a fallback for precision, and another enhancing repair tools with insights from LLMs. Both approaches will be evaluated for effectiveness.
@@ IV-B4 Proof Updater
The proposed proof updater will keep unit proofs synchronized with software changes. When the software changes, the updater will identify affected unit proofs, assess their validity, and apply necessary updates to maintain alignment with the software. To ground the updater’s design empirically, we will study how software changes impact unit proofs, what invalidates them, and how they are fixed.
To validate unit proofs after a change, the updater will reverify affected functions and compare the new results with prior verification data, using metrics such as coverage and violation reports based on the empirical study. Initially, the proof repairer (section IV-B3) will be used to resolve invalidated proofs. In future iterations, we plan to develop an LLM agent that, informed by empirical insights, can propose specific updates to the proof that realigns it with the changed software.
@@ VI Conclusion
Data availability: Our anonymous artifact contains the data for section III. See https://github.com/anon-dim98/icse25-nier.

### v2

@@ Abstract
Software underlies many critical digital and cyber-physical systems. Code-level defects, such as arithmetic errors and memory corruption vulnerabilities, compromise system safety and security and cause significant losses [1]. For example, a null-pointer bug in CrowdStrike Falcon in 2024 impacted global transportation systems and caused over $5 billion in damages [2]. Traditional engineering practices, such as testing [3, 4], program analysis [5], and runtime sandboxing[6], can detect or mitigate such defects. However, these methods validate software only approximately and cannot guarantee correct behavior on all inputs. In contrast, formal verification [7] proves software correctness, guaranteeing no code-level defects. This raises an important (and perennial) question [8]: how to make formal verification more cost-effective for software engineers?
Engineering tools and processes help reduce the cost of applying a technology. For formal methods, there are code-level verification tools, e.g., the C and Rust bounded model checkers (CBMC[9], Kani [10]). To improve both engineering and tool scalability, engineers decompose the software and verify individual functions within it via unit proofs [11]. These unit proofs model the function’s interaction with other software components and verify the function using this model. We refer to the process of developing, using, and maintaining unit proofs as unit proofing. Organizations such as AWS [12] and ARM [13] have adopted unit proofing to ensure software correctness and memory safety.
@@ II-A Formally Verifying Software Implementations
Many tools verify software implementations using a variety of formal methods [15, 16, 9]. This paper focuses on tools that use Bounded Model Checking (BMC) [17] as they require the least human effort [18], can automatically verify properties like memory safety, and are being adopted by major software organizations [12, 13]. We refer the reader to other works for comparisons of such tools [19, 18].
@@
BMC has been applied in both industry (e.g., AWS libraries [12], ARM confidential computing firmware [13]) and open-source projects (e.g., Firecracker, Hifitime, S2N-Quic [21]). Beyer et al. [22] report that BMC tools find significantly more bugs in less time compared to testing and fuzzing tools. However, BMC struggles with state space explosions when applied to large software systems.
@@ II-B Unit Proofing for Enabling Code-level Verification
To apply BMC tools in practice, several papers have described similar engineering processes [12, 13]. One example is AWS’s verification of embedded software and data structure libraries [12]. In their process, engineers develop unit proofs and use them to verify individual functions in the software. We refer to this process as “Unit Proofing”.
Unit proofing comprise four stages. (1) Software is first decomposed into manageable units. (2) Unit proofs are then created to configure verification and model the unit’s environment, including input, resources, and undefined functions (Figure 1). (3) Initial proofs may include misconfigurations, inaccurate input or function models, causing prolonged verification, low coverage, or false positives and negatives. These faults are investigated and the proofs repaired. (4) Finally, as software evolves, proofs must be updated to remain aligned with their respective units. Manually completing these steps, as reported by Chong et al. [12], is costly and error-prone.
Unit proofing shares similarities with unit testing and compositional verification. Like unit tests, AWS’s unit proofs are developed together with the software, stored in the same repository, and executed in continuous integration workflows [12]. However, unit tests check correctness on specific inputs, while unit proofs verify correctness across all possible inputs within the model [23]. Compared with compositional verification, unit proofing is an engineering process that enables compositional bounded model checking. There are works that automate compositional bounded model checking on real software [24, 25] and abstract software models [26, 27]. We focus on unit proofing due to its adoption by major software companies, including AWS and ARM.
@@ III Exploratory Study of Unit Proofing
We evaluated this question using FreeRTOS unit proofs and reported vulnerabilities (CVEs). FreeRTOS was chosen due to its widespread use and available unit proofs [29]. CVEs were used as they typically reflect critical memory safety vulnerabilities, which the unit proofs intend to address [29].
@@ III-A Methodology
Vulnerability Recreation: Following the vulnerability description, we removed any input validations introduced in the function that fixed the vulnerability. Attempts to build and evaluate versions that predate the vulnerability fix failed, mostly due to unavailable or incompatible dependencies.
@@ III-B Results
Only 5/11 recreated vulnerabilities were exposed by FreeRTOS unit proofs (Table I). With changes to the unit proofs, all vulnerabilities were exposed and one new high-severity vulnerability found (CVE-2024-38373). Unit proofs are effective, but unit proof engineers can err.
@@
| CVE-2018-16527 | No | Insuff. Input Model | Development |
| CVE-2018-16600 | No | Insuff. Input Model | Development |
| CVE-2018-16603 | No | Insuff. Input Model | Development |
@@ III-C1 Insufficient Verification Properties (CVE-2018-16523)
BMC tools can verify program-agnostic properties, but require the properties to be enabled. The unit proof for CVE-2018-16523, a division-by-zero defect, missed the issue because the division-by-zero property was disabled 111From CBMC v6, this property is enabled by default.. The rationale behind disabling specific correctness checks remains unclear.
@@
Unit proofs must model inputs broadly enough to cover all call sites yet precisely enough to avoid false positives. CVE-2018-16527 only occurs when the header of the input packet is truncated. The unit proof for the affected function incorrectly assumed the input packet always contained the header.
@@ III-C4 Out-of-date Unit Proofs (CVE-2018-16525)
The vulnerable function’s signature was changed but the new signature was not propagated to the unit proof. As CBMC supported making calls to undefined functions and only reported coverage on the reachable lines of code, the updated function was no longer verified. When we fixed this unit proof, we exposed and reported a new vulnerability 222In response to our disclosure, CBMC’s default behavior for undefined functions was changed to report errors instead of warnings. See https://github.com/diffblue/cbmc/pull/8292. .
@@ IV Research Agenda: A Unit Proofing Framework
IV-A Unit Proofing Stages and Open Questions
Here, we discuss the unit proofing stages introduced in section II-B and identify Open Questions (OQs) within them.
@@ IV-A Unit Proofing Stages and Open Questions
Unit proofing involves decomposing software into function units and verifying each one. The decomposition strategy impacts the number of unit proofs, the complexity of their models, and verification tractability. Smaller units are solvable but require creating more unit proofs to cover the software. Additionally, the effort to create a unit proof depend on the complexity of the models it requires. Units with complex input structures or that access multiple shared resources may require more effort to create unit proof models. Hence, an optimal decomposition should minimize unit proofing efforts but ensure all units are solvable.
@@
Software evolves, and section III shows that failing to update unit proofs leads to incorrect verification. Since unit proofs verify interacting functions, changes in one proof can invalidate others. Frequent software changes can make it impractical to regenerate or validate unit proofs after every change [12]. Understanding how software changes will affect unit proofs, and repairing misaligned unit proofs, will ensure unit proofs remain consistent with the software.
@@ IV-B1 Software Decomposer
This tool will employ novel algorithms to minimize unit count while ensuring verification completes within a set time. This requires new methods to estimate model complexity and verification time. Building on research linking program features to specification size and effort [32, 33], we will analyze unit proofs to understand how program features influence complexity and duration. These insights will guide a cost estimator design that evaluates function features to estimate complexity and verification time and a decomposer algorithm to iteratively replace the most expensive functions with models. We will also explore a bottom-up approach using graph decomposition algorithms [34, 35] to partition call graphs, balancing verification time and modeling complexity to minimize overall costs.
@@ IV-B2 Proof Builder
A unit proof requires sound and complete models for accurate verification. In stateful systems or those using shared resources, the resource model also depends on prior execution sequences. Traditional learning-based assume-guarantee reasoning [36, 37] refines assumptions iteratively using counterexamples, but applying this to functions with multiple interactions is costly due to the need for repeated model checking [38]. To address this, we propose to use a property-violation guided learning approach, starting with generic proof models and refining the models only until all reachable code is covered and reported violations are resolved. Refinements will be informed by interventions from the proof repairer (section IV-B3). The learned models will then be propagated as assertions to the respective components, with violations indicating defects. This approach will also help indicate specific execution paths that trigger a defect.
@@ IV-B3 Proof Repairer
The proof repairer will detect faults in unit proofs and intelligently suggest fixes. To design the repairer, we will develop a taxonomy of unit proof faults and interventions by empirically studying existing and new unit proofs, assessing their coverage, violation reports, and bug detection abilities, and identifying necessary interventions for different faults. Investigating and fixing proof faults requires a deep understanding of program semantics. Hence, using the developed taxonomy, the repairer will detect faults by combining the fast but imprecise program analysis techniques the code reasoning abilities of large language models (LLMs) [39, 40]. We will explore two hybrid approaches: one using search-based repairs with LLMs as a fallback for precision, and another enhancing repair tools with insights from LLMs. Both approaches will be evaluated for effectiveness.
@@ IV-B4 Proof Updater
The proposed proof updater will keep unit proofs synchronized with software changes. When the software changes, the updater will identify affected unit proofs, assess their validity, and apply necessary updates. An empirical study will examine how software changes impact proofs, what invalidates them, and how they are fixed. To validate updates, the updater will reverify affected functions, comparing results with prior data using metrics like coverage and violation reports. Initially, the proof repairer (section IV-B3) will be used to resolve invalidated proofs. In future iterations, we plan to develop an LLM agent that, informed by empirical insights, can propose specific updates to the proof that realigns it with the changed software.
@@ VI Conclusion
Acknowledgement: This project is being supported by a Rolls Royce research grant.
Data availability: Our artifact contains the data for section III. See https://doi.org/10.5281/zenodo.14660073.